OpenStudy (anonymous):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth?

OpenStudy (jojokiw3):

31 - 19 = 12. 198 - 5 = 3. 12/ 3 = 4. Each multiple choice question is worth 4.

OpenStudy (anonymous):

omg thank you so much :D

OpenStudy (anonymous):

Umm i have another question.

OpenStudy (anonymous):

In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March?

OpenStudy (anonymous):

I think it is $2940 that they had in march

OpenStudy (anonymous):

is it right?

OpenStudy (jojokiw3):

No. Now think this. We have the equation 15% = $210 dollars. So, 15% = 210. You want to figure out how much money 5% is, so there are three 5%s in 15%. Divide 210 by 3, which is $70 dollars. Now how many 5%s are there in a 100%? Can you figure it out from here?